OpenStudy (anonymous):

how do I integrate the following: (x^2)*sqrt(4x^2 +1)

OpenStudy (anonymous):

put everything in the square root to be square of a new variable say sqrt(4x^2 +1)=t or (4x^2 +1)=t^2..